OpenStudy (anonymous):

3x+y=6 4x+2y=8 using substitution to solve each equations.

OpenStudy (bahrom7893):

y=6-3x 4x+2(6-3x)=8 y=6-3x 4x+12-6x=8 etc.

OpenStudy (anonymous):

what else do i do

OpenStudy (anonymous):

you simplify. 4x + 12 - 6x = 8 10x = -4 x = -2/5 plug that back into y = 6 - 3x y = 6 - 3(-2/5) y = 4.8 answer is (-0.4 , +4.8) == (-2/5, +24/5)

OpenStudy (anonymous):

now really

OpenStudy (anonymous):

that how you got setup idk